# Asymptotic behaviors and dynamics of degenerate and mixed solitons for the coupled Hirota system with strong coherent coupling effects

### v1

@@ D. Case 4: $l_{13}^{2}+l_{14}^{2}=0$ and $l_{33}^{2}+l_{34}^{2}=0$
$\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}-}\\ q_{2}^{{\rm line}-}\end{array}\right)$ (243)
$={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\end{array}\right).$ (246)
@@
$\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}+}\\ q_{2}^{{\rm line}+}\end{array}\right)={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{m_{11}-\frac{l_{13}}{l_{14}}m_{12}}{\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{{\rho}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{{\rho}}{2}\right)\end{array}\right).$ (274)
@@ 5. Interactions of the mixed solitons
  [Uncaptioned image]   [Uncaptioned image]
@@
Thus, in the case of $l_{13}^{2}+l_{14}^{2}=0$ and $l_{33}^{2}+l_{34}^{2}\neq 0$ $\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\neq 0\right)$, the interaction is elastic for the degenerate soliton consisting of $S^{1}$, $S^{2}$ and $S^{\rm line}$ in both $q_{1}$ and $q_{2}$ components, but inelastic for the bell-shaped soliton $S^{3}$, as illustrated in Figs. 11(a1-a2) and 11(b1-b2). Figs. 11(a1-a2) display that the bell-shaped soliton $S^{3}$ transitions from a slightly higher single-hump profile to a lower double-hump profile in both $q_{1}$ and $q_{2}$ components. Figs. 11(b1-b2) reveal that in the $q_{1}$ component, the soliton changes from a notably higher single-hump to a lower double-hump profile, while in the $q_{2}$ component, it shifts from an invisible hump to a double-hump profile, indicating a wave energy exchange between the $q_{1}$ and $q_{2}$ components. Comparing Figs. 11(a1-a2) with Figs. 11(c1-c2), we observe that the line soliton in the degenerate soliton $S^{\rm line}$ disappears in both $q_{1}$ and $q_{2}$ components when $m_{11}-\frac{l_{13}}{l_{14}}m_{12}=0$.
@@
  [Uncaptioned image]   [Uncaptioned image]
@@
In this context, it is noted that the intensity redistribution between soliton components cannot be conclusively determined from asymptotic amplitudes alone. Degenerate soliton branches may exhibit some coherent interaction phenomena during a longer interaction region, as illustrated in Figs. 12(a1-a2) and 12(b1-b2). With $m_{11}-\frac{l_{13}}{l_{14}}m_{12}\neq 0$, Figs. 12(a1-a2) and 12(b1-b2) reveal that degenerate soliton branches $S^{1}$ and $S^{\rm line}$ experience coherent interaction before interacting with $S^{3}$ in the $q_{1}$ and $q_{2}$ components. However, when $m_{11}-\frac{l_{13}}{l_{14}}m_{12}=0$, the line soliton in the degenerate soliton $S^{\rm line}$ disappears, resulting in an elastic interaction without coherent effects in both $q_{1}$ and $q_{2}$ components, as depicted in Figs. 12(c1-c2).
@@
  [Uncaptioned image]   [Uncaptioned image]
@@ 6. Strong coherence phenomena and robust analysis
Based on Solutions (17), we can know that parameter $\varepsilon$ is independent of the initial intensities $|q_{1}(x,0)|$ and $|q_{2}(x,0)|$ of degenerate solitons, which are shown in Figs. 13(a1-a2). Figs. 13(b1-b2) show the intensity profiles of degenerate solitons at $t=2$. We can see that the wave energy increases significantly in the $q_{1}$ component, but decays significantly in the $q_{2}$ component. Such phenomena can only occur in a coherent systems, which are different from the solitons in some incoherent systems, e.g., the conventional coupled NLS system or coupled Hirota system. More importantly, when we increase $\varepsilon$, the wave energy changes more sharp. Particularly, when we choose $\varepsilon=2$, the soliton intensities in the $q_{1}$ component expand about $7$ times to maximum from the initial state, but decrease to only $1/7$ of their initial intensities in the $q_{2}$ component, see the blue line in Figs. 13(b1-b2). Figs. 14 show the strong coherence degenerate solitons. In other words, the higher-order perturbation parameter $\varepsilon$ distinctly affects the coherence of degenerate solitons. The larger $\varepsilon$ leads to the stronger coherence.
@@
$(|q_{1}|^{2})_{t}={\rm i}(q_{1,x}q_{1}^{*}-q_{1,x}^{*}q_{1})_{x}-2{\rm i}(q_{1}^{*2}q_{2}^{2}-q_{1}^{2}q_{2}^{*2})-\varepsilon(q_{1,xx}q_{1}^{*}+q_{1,xx}^{*}q_{1}-|q_{1,x}|^{2})_{x}$
$-6\varepsilon\left[(|q_{1}|^{2}+|q_{2}|^{2})|q_{1}|^{2}-\frac{1}{2}|q_{1}|^{4}\right]_{x}+3\varepsilon\left[q_{1}^{*2}(q_{2}^{2})_{x}+q_{1}^{2}(q_{2}^{*2})_{x}\right],$ (290a)
$(|q_{2}|^{2})_{t}={\rm i}(q_{2,x}q_{2}^{*}-q_{2,x}^{*}q_{2})_{x}+2{\rm i}(q_{1}^{*2}q_{2}^{2}-q_{1}^{2}q_{2}^{*2})-\varepsilon(q_{2,xx}q_{2}^{*}+q_{2,xx}^{*}q_{2}-|q_{2,x}|^{2})_{x}$
$-6\varepsilon\left[(|q_{1}|^{2}+|q_{2}|^{2})|q_{2}|^{2}-\frac{1}{2}|q_{2}|^{4}\right]_{x}+3\varepsilon\left[q_{2}^{*2}(q_{1}^{2})_{x}+q_{2}^{2}(q_{1}^{*2})_{x}\right],$ (290b)
@@
Thus the coherent coupling terms of System (1) actually have two parts, and the higher-order perturbation parameter $\varepsilon$ play a significant role on the coherence effect. The energy exchange between $q_{1}$ and $q_{2}$ components may become stronger when $|\varepsilon|$ increases. It is noticed that such phenomena of System (1) are completely different from the matrix NLS system and the vector Hirota system without coherent coupling [5, 35, 36].
@@
Finally, we numerically analyze the relationship between the robustness of such solitons and $\varepsilon$. Here, a small rapidly decaying perturbation to the initial condition is considered, i.e., $q_{1,2}^{\delta}(x,t=0)=q_{1,2}(x,0)+\delta f(x)$, where $\delta$ is a small real constant and $f(x)=e^{-x^{2}}$. For a fixed small perturbation and different $\varepsilon$, the initial intensities $q_{1}$ and $q_{2}$ are almost the same. We can see that the soliton evolution keep basically stable in a short time $t=4$ for $\varepsilon=\frac{1}{25}$, as seen in Figs. 15(a1-b1). However, the robustness of such degenerate solitons become weaker with increasing $\varepsilon$, as displayed in Figs. 15(a2-b2) and (a3-b3). The coherence effects have a negative effect on the dynamics stability of the degenerate solitons. It is noted that such the degenerate solitons are generally dynamic unstable under a perturbation in a large time.
@@ 7. Conclusions
In this paper, we have investigated the asymptotic behaviors and dynamics of degenerate and mixed solitons for the coupled Hirota system (1), which can describe the optical pulse propagation in isotropic nonlinear medium. Firstly, using Binary DT (6), we have derived Solutions (17) to represent the degenerate soliton solutions with two eigenvalues that are conjugate to each other. We have obtained three types of degenerate solitons, with their asymptotic expressions detailed in Expressions (36) and (51). Notably, these degenerate solitons exhibit time-dependent velocities, as illustrated in Expressions (39) and (54). Figs. 1 have depicted the asymptotic soliton branch along the straight line $\theta=\mathcal{O}(1)$, while Figs. 2-5 have displayed those three types of the degenerate solitons. Furthermore, the relative distance between the two asymptotic solitons, $S^{1}$ and $S^{2}$, has been observed to increase logarithmically with the absolute value of the higher-order perturbation parameter $\varepsilon$. Figs. 6-8 have compared Asymptotic Expressions (55) and (56) with Solutions (24), demonstrating that the asymptotic solitons are generally in good agreement with the degenerate solitons.
@@
Secondly, we have asymptotically and graphically revealed four interaction mechanisms between a degenerate soliton and a bell-shaped soliton, as shown in Figs. 9-12: Figs. 9 have displayed an elastic interaction with a few position shifts at the interaction regions; Figs. 10 have shown the inelastic interaction for the degenerate soliton but elastic interaction for the bell-shaped soliton; Figs. 11 have illustrated the elastic interaction for the degenerate soliton and inelastic interaction for the bell-shaped soliton; Figs. 12 have exhibited the coherent interaction during a longer interaction region, where $m_{11}-\frac{l_{13}}{l_{14}}m_{12}\neq 0$, and elastic interaction, where $m_{11}-\frac{l_{13}}{l_{14}}m_{12}=0$.
Finally, we have uncovered and discussed a special degenerate vector soliton which shows significant coherence effects in Figs. 3, and analyzed the relationship between the robustness of such solitons and higher-order perturbation parameter $\varepsilon$ through the numerical evolution. Figs. 13 have illustrated that the wave energy between two components will exchange significantly when $|\varepsilon|$ increases. Figs. 14 have shown a strong coherence degenerate vector soliton which has remarkable intensities change around time $t=0$. Figs. 15 have exhibited that the robustness of such degenerate solitons become weaker with increasing $|\varepsilon|$.
@@ Acknowledgments
This work has been supported by the National Natural Science Foundation of China under Grant Nos. 12305001 and 12205029, by the Fundamental Research Funds for the Central Universities (Grant Nos. 2024CDJXY018 and 2023MS163), and by the Natural Science Foundation of Hebei Province under Grant No. A2021502003.

### v2

@@ D. Case 4: $l_{13}^{2}+l_{14}^{2}=0$ and $l_{33}^{2}+l_{34}^{2}=0$
$\hskip-48.36958pt\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}-}\\ q_{2}^{{\rm line}-}\end{array}\right)$ (243)
$\hskip-8.5359pt={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\end{array}\right).$ (246)
@@
$\hskip-48.36958pt\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}+}\\ q_{2}^{{\rm line}+}\end{array}\right)={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{m_{11}-\frac{l_{13}}{l_{14}}m_{12}}{\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{{\rho}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{{\rho}}{2}\right)\end{array}\right).$ (274)
@@ 5. Interactions of the mixed solitons
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@
Thus, in the case of $l_{13}^{2}+l_{14}^{2}=0$ and $l_{33}^{2}+l_{34}^{2}\neq 0$ $\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\neq 0\right)$, the interaction is elastic for the degenerate soliton consisting of $S^{1}$, $S^{2}$ and $S^{\rm line}$ in both $q_{1}$ and $q_{2}$ components, but inelastic for the bell-shaped soliton $S^{3}$, as illustrated in Figs. 10(a1-a2) and 10(b1-b2). Figs. 10(a1-a2) display that the bell-shaped soliton $S^{3}$ transitions from a slightly higher single-hump profile to a lower double-hump profile in both $q_{1}$ and $q_{2}$ components. Figs. 10(b1-b2) reveal that in the $q_{1}$ component, the soliton changes from a notably higher single-hump to a lower double-hump profile, while in the $q_{2}$ component, it shifts from an invisible hump to a double-hump profile, indicating a wave energy exchange between the $q_{1}$ and $q_{2}$ components. Comparing Figs. 10(a1-a2) with Figs. 10(c1-c2), we observe that the line soliton in the degenerate soliton $S^{\rm line}$ disappears in both $q_{1}$ and $q_{2}$ components when $m_{11}-\frac{l_{13}}{l_{14}}m_{12}=0$.
@@
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@
In this context, it is noted that the intensity redistribution between soliton components cannot be conclusively determined from asymptotic amplitudes alone. Degenerate soliton branches may exhibit some coherent interaction phenomena during a longer interaction region, as illustrated in Figs. 11(a1-a2) and 11(b1-b2). With $m_{11}-\frac{l_{13}}{l_{14}}m_{12}\neq 0$, Figs. 11(a1-a2) and 11(b1-b2) reveal that degenerate soliton branches $S^{1}$ and $S^{\rm line}$ experience coherent interaction before interacting with $S^{3}$ in the $q_{1}$ and $q_{2}$ components. However, when $m_{11}-\frac{l_{13}}{l_{14}}m_{12}=0$, the line soliton in the degenerate soliton $S^{\rm line}$ disappears, resulting in an elastic interaction without coherent effects in both $q_{1}$ and $q_{2}$ components, as depicted in Figs. 11(c1-c2).
@@
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@ 6. Strong coherence phenomena and robust analysis
Based on Solutions (17), we can know that parameter $\varepsilon$ is independent of the initial intensities $|q_{1}(x,0)|$ and $|q_{2}(x,0)|$ of degenerate solitons, which are shown in Figs. 13(a1-a2). Figs. 13(b1-b2) show the intensity profiles of degenerate solitons at $t=2$. We can see that the wave energy increases significantly in the $q_{1}$ component, but decays significantly in the $q_{2}$ component. Such phenomena can only occur in a coherent systems, which are different from the solitons in some incoherent systems, e.g., the conventional coupled NLS system or coupled Hirota system. More importantly, when we increase $\varepsilon$, the wave energy changes more sharp. Particularly, when we choose $\varepsilon=2$, the soliton intensities in the $q_{1}$ component expand about 777 times to maximum from the initial state, but decrease to only 1/71/71 / 7 of their initial intensities in the $q_{2}$ component, see the blue line in Figs. 12(b1-b2). Figs. 13 show the strong coherence degenerate solitons. In other words, the higher-order perturbation parameter $\varepsilon$ distinctly affects the coherence of degenerate solitons. The larger $\varepsilon$ leads to the stronger coherence.
@@
$\hskip-42.67912pt(|q_{1}|^{2})_{t}={\rm i}(q_{1,x}q_{1}^{*}-q_{1,x}^{*}q_{1})_{x}-2{\rm i}(q_{1}^{*2}q_{2}^{2}-q_{1}^{2}q_{2}^{*2})-\varepsilon(q_{1,xx}q_{1}^{*}+q_{1,xx}^{*}q_{1}-|q_{1,x}|^{2})_{x}$
$\hskip-28.45274pt-6\varepsilon\left[(|q_{1}|^{2}+|q_{2}|^{2})|q_{1}|^{2}-\frac{1}{2}|q_{1}|^{4}\right]_{x}+3\varepsilon\left[q_{1}^{*2}(q_{2}^{2})_{x}+q_{1}^{2}(q_{2}^{*2})_{x}\right],$ (290a)
$\hskip-42.67912pt(|q_{2}|^{2})_{t}={\rm i}(q_{2,x}q_{2}^{*}-q_{2,x}^{*}q_{2})_{x}+2{\rm i}(q_{1}^{*2}q_{2}^{2}-q_{1}^{2}q_{2}^{*2})-\varepsilon(q_{2,xx}q_{2}^{*}+q_{2,xx}^{*}q_{2}-|q_{2,x}|^{2})_{x}$
$\hskip-28.45274pt-6\varepsilon\left[(|q_{1}|^{2}+|q_{2}|^{2})|q_{2}|^{2}-\frac{1}{2}|q_{2}|^{4}\right]_{x}+3\varepsilon\left[q_{2}^{*2}(q_{1}^{2})_{x}+q_{2}^{2}(q_{1}^{*2})_{x}\right],$ (290b)
@@
Thus the coherent coupling terms of System (1) actually have two parts, and the higher-order perturbation parameter $\varepsilon$ play a significant role on the coherence effect. The energy exchange between $q_{1}$ and $q_{2}$ components may become stronger when $|\varepsilon|$ increases. Such phenomena of System (1) are completely different from the matrix NLS system and the vector Hirota system without coherent coupling [8, 43, 44].
@@
Finally, we numerically analyze the relationship between the robustness of such solitons and $\varepsilon$. Here, a small rapidly decaying perturbation to the initial condition is considered, i.e., $q_{1,2}^{\delta}(x,t=0)=q_{1,2}(x,0)+\delta f(x)$, where $\delta$ is a small real constant and $f(x)=e^{-x^{2}}$. Here, to obtain a high resolution, we use the Fourier pseudospectral discretization with at least $N=2^{9}$ Fourier modes to deal with variable $x$ and the fourth-order Runge-Kutta scheme with a step size less than $10^{-4}$ is used to the time discretization. The computation interval is taken as [15,15][-15,15][ - 15 , 15 ], which is large enough to ignore the effect of periodic boundary errors. The corresponding simulations for each of the cases presented only took a few minutes of computer time on a standard desktop computer.
For a fixed small perturbation and different $\varepsilon$, the initial intensities $q_{1}$ and $q_{2}$ are almost the same. We can see that the soliton evolution keep basically stable in a short time $t=4$ for $\varepsilon=\frac{1}{25}$, as seen in Figs. 14(a1-b1). However, the robustness of such degenerate solitons become weaker with increasing $\varepsilon$, as displayed in Figs. 14(a2-b2) and (a3-b3). The coherence effects have a negative effect on the dynamics stability of the degenerate solitons. It is noted that such the degenerate solitons are generally dynamic unstable under a perturbation in a large time.
@@ 7. Conclusions and discussions
In this paper, we have investigated the asymptotic behaviors and dynamics of degenerate and mixed solitons for the coupled Hirota system (1), which can describe the optical pulse propagation in isotropic nonlinear medium. Firstly, using Binary DT (6), we have derived Solutions (17) to represent the degenerate soliton solutions with two eigenvalues that are conjugate to each other. We have obtained three types of degenerate solitons, with their asymptotic expressions detailed in Expressions (36) and (51). Notably, these degenerate solitons exhibit time-dependent velocities, as illustrated in Expressions (39) and (54). Based on those solutons and expressions, we have obtained the following prominent results:
  • We have revealed that higher-order perturbation parameter $\varepsilon$ has a significant impact on the coherence and robustness of solitons. Specifically, the relative distance between asymptotic solitons increases logarithmically with the increase of the high-order perturbation parameter $|\varepsilon|$.
@@
  • We have asymptotically and graphically discovered four different interaction mechanisms between degenerate solitons and bell-shaped solitons, including elastic interactions with position shifts, interactions where degenerate solitons are inelastic but bell-shaped solitons are elastic, interactions where degenerate solitons are elastic but bell-shaped solitons are inelastic, and coherent interactions that occur over longer interaction regions.
  • By means of numerical simulation, we have shown that the coherence of the degenerate soliton is strongly affected by the parameter $\varepsilon$, and its robustness decreases with the increase of $|\varepsilon|$. This discovery highlights the importance of considering higher-order effects in practical applications.
Besides, the findings in this paper have several physical implications, which are described in detail as follows: (1) Detailed analysis of degenerate solitons and mixed solitons provides new insights into the behavior of optical pulses in isotropic nonlinear media. (2) The research results show that the high-order effect significantly affects the coherence and robustness of solitons. (3) The research results on energy redistribution and coherence effects indicate that these solitons have potential application value in fields such as nonlinear optics and fluid mechanics.
Moreover, although the binary Darboux transformation method adopted in this paper successfully derive degenerate and mixed soliton solutions for the coupled Hirota system with strong coherent coupling effects, it also has some limitations. For instance, its solutions are sensitive to initial conditions and parameter choices, and the calculation process is complex, which may introduce numerical instability. Although theoretical analysis and numerical simulations provide important insights into soliton dynamics, the lack of direct experimental verification and the potential impact of higher-order perturbation parameters on the robustness and stability of solitons in practical applications are issues. Therefore, future research needs to further explore these limitations and verify the applicability of the solutions in a wider range of systems.
While the primary focus of our study is on the coupled Hirota system in the context of nonlinear optics, the methodology and findings have broader implications, particularly in the field of fluid mechanics, such as studying the wave dynamics in fluids, turbulence phenomena, and designing fluid systems through the analysis of soliton propagation and interaction. The higher-order effects considered in our model, such as third-order dispersion and self-steepening, are particularly important for accurately describing the propagation of ultrashort pulses in the marine environment. Our research can enhance the understanding of these phenomena and their potential impacts on ocean dynamics, coastal processes and fluid-structure interactions such as the response of offshore structures to extreme wave events. Meanwhile, future research directions and scope should include: (1) Experimentally verifying theoretical results; (2) Extending the method to other fluid systems; (3) Using numerical simulations to study more complex fluid environments; (4) Studying practical applications in engineering and technology, such as designing fluid devices and wave energy converters. These efforts will help to realize the potential application of this research in fluid mechanics.
@@ Acknowledgments
This work has been supported by the National Natural Science Foundation of China under Grant Nos. 12305001 and 12205029, by the Fundamental Research Funds for the Central Universities (Grant Nos. 2024CDJXY018 and 2025MS183), and by the China Scholarship Council (No. 202306050162).